The mass (in kg), position coordinates (in m) and velocity (in m s−1
) of a system of
three particles A, B and C are as follows:
Particle Mass Position Velocity
A 2.0 kg (0, 0) 0
r
B 1.0 kg (2, 0) 4.0 j
ˆ
C 3.0 kg (1, 1) 2.0 i
ˆ
Calculate the coordinates of the centre of mass of the system and its momentum.
1
Expert's answer
2019-10-01T12:47:08-0400
"\bold{r_m}=\frac{(0,0)2+(2,0)+(1,1)3}{2+1+3}=(0.83,0.5)\ m"

"\bold{P}=1(0,4)+3(2,0)=(6,4)\frac{kg\cdot m}{s}"